# Simple Guidance Mechanisms for Discrete Diffusion Models

### v1

@@ 4.2 Improved Likelihood Bounds in Continuous Time
We now present our contribution to uniform noise discrete diffusion modeling. We leverage the above formulation to develop an improved ELBO by taking $T\rightarrow\infty$ and analyzing each term ${\mathcal{L}_{\text{recons}}},{\mathcal{L}_{\text{diffusion}}},{\mathcal{L}_{\text{prior}}}$ in (13). This yields three improvements: (1) a simple and elegant closed-form expression for the ELBO that is easier to reason about; (2) an analytical reduction of ${\mathcal{L}_{\text{recons}}},{\mathcal{L}_{\text{prior}}}$ to zero, which tightens the ELBO; (3) a further tightening via the continuous-time extension of ${\mathcal{L}_{\text{diffusion}}}$, as in Kingma et al. (2021) and Sahoo et al. (2024a; b).
@@
Reconstruction Loss (${\mathcal{L}_{\text{recons}}}$)  For ${\mathcal{L}_{\text{recons}}}$, if our noise schedule is such that $T\rightarrow\infty\implies\alpha_{t(0)}=1$ (i.e., $t(0)=1/T$), then the marginal $q({\mathbf{z}}_{\frac{1}{T}}\mid{\mathbf{x}}_{0})\rightarrow\mathrm{Cat}({\mathbf{z}}_{\frac{1}{T}};{\mathbf{x}}_{0})$. That is, in the limit, the first latent vector is identically equal to the clean data. We can thus parameterize our denoising network such that at time $t(0)$ the function simply copies its inputs: ${\mathbf{x}}_{\theta}({\mathbf{z}}_{\frac{1}{T}},1/T)={\mathbf{z}}_{\frac{1}{T}}.$ Additionally, we note that our choice of parameterization for $p_{\theta}$ implies that $p_{\theta}({\mathbf{x}}_{0}\mid{\mathbf{z}}_{{\frac{1}{T}}})={\mathbf{x}}_{\theta}({\mathbf{z}}_{{\frac{1}{T}}},1/T).$ Thus in the continuous time limit, we have:
$\lim_{T\rightarrow\infty}\mathbb{E}_{q}[{\mathcal{L}_{\text{recons}}}]=\lim_{T\rightarrow\infty}\mathbb{E}_{q}[\log p_{\theta}({\mathbf{x}}_{0}\mid{\mathbf{z}}_{\frac{1}{T}})]=\lim_{T\rightarrow\infty}\mathbb{E}_{q}[\log(\langle{\mathbf{x}}_{0},{\mathbf{x}}_{\theta}({\mathbf{z}}_{\frac{1}{T}},1/T)\rangle)]=0.$
Diffusion Loss (${\mathcal{L}_{\text{diffusion}}}$)  Turning finally to the diffusion loss term ${\mathcal{L}_{\text{diffusion}}}$, we first define the shorthand $D_{\mathrm{KL}}[q_{t}||p_{\theta}]=D_{\mathrm{KL}}[[q({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t},{\mathbf{x}}_{0})||p_{\theta}({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t})]]$ and then re-write this loss term as an expectation over $t$ uniformly sampled from $1/T,2/T,\ldots,T$:
@@
where ${\mathbf{x}}_{i}$ denotes the $i^{\text{th}}$ index of a vector ${\mathbf{x}}$, $\bar{\mathbf{x}}=N\alpha_{t}{\mathbf{x}}+(1-\alpha_{t}){\bm{1}}$, $\bar{\mathbf{x}}_{\theta}=N\alpha_{t}{\mathbf{x}}_{\theta}+(1-\alpha_{t}){\bm{1}}$, and we define $i=\arg\max_{j\in[N]}({\mathbf{z}}_{t})_{j}$ to be the non-zero entry of ${\mathbf{z}}_{t}$.
Combining our arguments regarding ${\mathcal{L}_{\text{prior}}}$ and ${\mathcal{L}_{\text{recons}}}$ with (17), yields our final tight variational bound:
@@ 5 Experiments
Datasets  For our language modeling experiments we examine several distinct discrete domains: reference genomes from ten diverse species (Species10), the small molecule dataset known as QM9 (Ruddigkeit et al., 2012; Ramakrishnan et al., 2014), where molecules are represented by a string of characters known as SMILES (Weininger, 1988), discretized images from CIFAR10 (Krizhevsky et al., 2009), and three NLP datasets consisting of text8 (Mahoney, 2011), Amazon Review Polarity (McAuley & Leskovec, 2013; Zhang et al., 2015), and the one billion words dataset (LM1B; Chelba et al. (2014)). These datasets cover a range of domains and vocabularies of varying sizes (see Table 1). For guidance experiments, we explore species-specific sequence generation, molecular property maximization, and class-conditional image generation.
@@ 5.1 Language Modeling with Uniform Noise Discrete Diffusion
Our language modeling experiments reveal that (1) contrary to a widely-held belief, uniform noise diffusion can attain state-of-the-art performance on small vocabulary datasets (Table 1), and that (2) our UDLM are state-of-the-art among uniform noise diffusion models (Tables 2 and 3).
Figure 2: Increasing $T$ to infinity improves language modeling. Validation perplexity (PPL $\downarrow$) on Amazon polarity dataset for uniform noise diffusion using finite $T$ vs. UDLM ($T=\infty$).
In Table 1, despite previous evidence indicating that absorbing-state discrete diffusion greatly outperforms uniform noise, we find a more nuanced story. Namely, for smaller vocabulary regimes, which are common for scientific applications, the gap between MDLM and UDLM is negligible, with UDLM even outperforming absorbing-state diffusion on certain datasets. Even within a single domain, we find a trend between vocabulary size and performance gap, with the text8 results of UDLM on par with MDLM, and a more persistent gap for the larger vocabulary experiments. Intuitively, these results follow from the observation that in larger vocabulary regimes, at every denoising step, uniform noise diffusion models need to predict over a combinatorially larger set of potential clean data sequences compared to absorbing-state; smaller vocabularies reduce this complexity.
Secondly, although we still find a persisting perplexity gap between absorbing and uniform noise discrete diffusion for larger vocabulary NLP datasets, we note that our uniform noise diffusion models trained with UDLM, closes this gap, as evidenced in Tables 2 and 3, where we show that UDLM attains the best reported uniform noise discrete diffusion language modeling performance.
Ablating Continuous Time Formulation  In Figure 2, we see the effect of increasing $T$ and more specifically using our continuous time ELBO. The curves represent validation perplexity on the Amazon Polarity dataset at various training steps, and we observe that increasing $T$ indeed improves language modeling with UDLM, i.e., $T=\infty$, performing best.
Table 1: UDLM performs best in smaller vocabulary regimes. Perplexity (PPL; $\downarrow$) on various datasets. Best values are bolded. The $**{}^{\text{*}}$ indicates values reported from early stopping on the validation set, since AR demonstrated overfitting; all other datasets reflect validation performance at the end of training. Taken from Sahoo et al. (2024a). \$Taken from Lou et al. (2023).
Vocab. Size AR MDLM UDLM
Bio. Species10 12 2.88 3.17 3.15
Chem. $**{}^{\text{*}}$ 40 2.19 2.12 2.02
Images CIFAR10 256 - 9.14 11.21
NLP text8 35 2.35\$ 2.62 2.71
@@
Table 2: UDLM outperforms previously reported uniform noise discrete diffusion models on text8 dataset. Values correspond to bits per character (BPC). Best values are bolded and best discrete uniform diffusion value is underlined. Taken from Lou et al. (2023). $**{}^{\text{*}}$aken from Shi et al. (2024)
@@
IAF/SCF [70] 1.88
AR Argmax Flow [21] 1.39
Discrete Flow [60] 1.23
@@
Mult. Diffusion [21] 1.72
MAC [53] 1.40
BFN [14] 1.41
D3PM Absorb [1] 1.45
SEDD Absorb [28] 1.39
MDLM [45] (Retrained) 1.38
$**{}^{\text{*}}$ [52] 1.37
GenMD4 [52] 1.34
@@
D3PM Uniform [1] 1.61
SEDD Uniform [28] 1.47
@@
Table 3: UDLM outperforms previously reported uniform noise discrete diffusion models on LM1B dataset. Values correspond to perplexity (PPL). Best values are bolded and best discrete uniform diffusion value is underlined. Taken from Sahoo et al. (2024a). $**{}^{\text{*}}$aken from Lou et al. (2023). \$Taken from Austin et al. (2021).
@@
BERT-Mouth [62] 142.89
D3PM Absorb [1] 77.50
Diffusion-LM [26] 118.62
DiffusionBert [62] 63.78
SEDD Absorb [28] 32.79
MDLM [45] 27.04
@@
D3PM ${}^{\text{\$}}$ [1] 137.9
SEDD $**{}^{\text{*}}$ [28] 40.25
@@ 5.2 Guided Discrete Diffusion
Our guidance results indicate that (1) classifier-free guidance is more useful when paired with diffusion models compared to AR (Table 4) and that (2) our proposed D-CBG is the best classifier-based method for discrete guidance, especially when combined with UDLM (Tables 5 and 6).
@@
Baselines  For guidance experiments, our primary baseline is the dominant approach for language modeling, AR. Our goal is to show that with similar, even slightly worse, language modeling performance, discrete diffusion models equipped with classifier-based and classifier-free guidance can outperform AR models on controlled generation. We compare to three flavors of guided AR models. The first is applying D-CFG to AR models. We also use the established control mechanisms of Plug-and-play language models (PPLM; Dathathri et al. (2019)) and FUDGE (Yang & Klein, 2021). To demonstrate the better performance of our D-CBG method, we compare to Gruver et al. (2024) (NOS) which can be viewed as an extension of PPLM to discrete diffusion models.
@@
Hyperparameters  For both our D-CFG and D-CBG, we vary the strength of the $\gamma$ parameter. Although the original FUDGE formulation simply uses $\gamma=1$, we also perform a search for this baseline. For PPLM and NOS, we vary the parameters of the Langevin sampling that is performed to update models’ hidden representation, namely the number of update steps $n$, the step size $\eta$, and the fluency KL weight $\gamma_{kl}$ (see Dathathri et al. (2019) and Gruver et al. (2024) for more details). For all methods, we display the best performing hyperparameter configuration in the main table results, deferring the full sweep results to Appendix D.
Species-specific Genome Generation  For genomic sequences, we evaluate D-CFG with diffusion models vs. with an AR model. We train models on sequences of length 32,768 nucleotides, using base-pair level tokenization and conditionally generate 64 sequences for each class. As quality measures for each species class, we compute the Jensen-Shannon (JS) distance between the $k$-mer frequencies of the generated sequences and those from the validation set, and report the mean JS across species (weighted by species frequency in the dataset), with smaller values indicating better $k$-mer distributional overlap between the ground truth and generated sequences (Sarkar et al., 2024). Additionally, we train a small classifier to distinguish between generated and validation set sequences and report the area under the receiver operator curve for this classifier (Disc. AUROC). Values closer to 0.5 indicate that the classifier is unable to distinguish between synthetic and true sequences (Sarkar et al., 2024). To measure the controllability, we train a separate classifier on the Species10 dataset and measure macro F1 score of this oracle classifier on the generated sequences.
Results of this experiment are presented in Table 4. For reference, we also provide metrics for randomly generating sequences with nucleotide frequencies proportional to species representation in the data. We find that both MDLM and UDLM are able to better generate sequences that match the desired control parameter, with higher F1 scores relative to AR. Moreover, UDLM is able to outperform MDLM in satisfying this control. Importantly, we find that only UDLM is amenable to increasing the guidance parameter $\gamma$, where its metrics improves while AR and MDLM metrics degrade. Finally, of note, the diffusion model generation for this experiment is accomplished with far fewer function evaluations compared to AR. Whereas AR must decode each of the 32,768 tokens, because MDLM and UDLM can decode multiple tokens in parallel, we generate with $T=512.$
@@
Table 4: Diffusion decoding with D-CFG is more controllable than AR for genomic sequences. Mean $\pm$ standard deviation reported from repeated generation of 640 sequences (64 per species) using five different random seeds. Best values are bolded.
Model Guidance 3-mer JS ($\downarrow$) 6-mer JS ($\downarrow$) Disc. AUROC ($\downarrow$) F1 ($\uparrow$)
Random - 0.13 0.22 1.00 0.07
AR D-CFGγ=1 0.03 $\pm$ 0.002 0.07 $\pm$ 0.004 0.53 $\pm$ 0.075 0.87 $\pm$ 0.011
AR D-CFGγ=2 0.05 $\pm$ 0.003 0.12 $\pm$ 0.004 0.9 $\pm$ 0.051 0.81 $\pm$ 0.009
AR D-CFGγ=3 0.07 $\pm$ 0.002 0.15 $\pm$ 0.002 0.97 $\pm$ 0.018 0.74 $\pm$ 0.022
MDLM D-CFGγ=1 0.02 $\pm$ 0.001 0.06 $\pm$ 0.001 0.51 $\pm$ 0.059 0.88 $\pm$ 0.011
MDLM D-CFGγ=2 0.05 $\pm$ 0.001 0.11 $\pm$ 0.003 0.74 $\pm$ 0.037 0.91 $\pm$ 0.009
MDLM D-CFGγ=3 0.11 $\pm$ 0.004 0.2 $\pm$ 0.005 0.93 $\pm$ 0.031 0.78 $\pm$ 0.007
UDLM D-CFGγ=1 0.02 $\pm$ 0.002 0.06 $\pm$ 0.003 0.52 $\pm$ 0.044 0.91 $\pm$ 0.01
UDLM D-CFGγ=2 0.05 $\pm$ 0.002 0.13 $\pm$ 0.002 0.61 $\pm$ 0.043 0.93 $\pm$ 0.009
UDLM D-CFGγ=3 0.08 $\pm$ 0.002 0.2 $\pm$ 0.003 0.87 $\pm$ 0.051 0.94 $\pm$ 0.007
Molecular Property Maximization  For the QM9 dataset, we investigate novel generation of sequences that maximize some property, either drug-likeness (QED; Bickerton et al. (2012)), see Table 5, or a count of the number of rings present in the molecule, see Table 6. We only report values for which we generated at least 50 novel sequences (out of 1,024). For D-CFG, we see that all methods perform comparably when maximizing drug likeness, but that MDLM and UDLM are better suited for the structural property of ring count. For classifier-based guidance mechanisms, we again find that both QED and ring count, D-CBG  with discrete diffusion best trades-off maximizing these properties while generating valid and novel sequences compared to AR or diffusion with NOS.
@@
Table 5: Guidance with discrete diffusion models better balances the generation of valid and novel molecules with maximizing the property of interest, drug likeness (QED), compared to AR. Validity, novelty, and mean QED for novel sequences are measured for generated sequences from each method. Mean $\pm$ standard deviation reported from repeated generation of 1,024 sequences using five different random seeds. Best mean values are bolded. “no approx.” subscript for D-CBG implies first order approximation was not used.
Method Guidance Valid ($\uparrow$) Novel ($\uparrow$) QED Mean ($\uparrow$)
Original Data - 133k 133k 0.47
@@
AR D-CFGγ=3 946.4 $\pm$ 8.99 79.4 $\pm$ 6.35 0.60 $\pm$ 0.00
MDLM D-CFGγ=3 317.4 $\pm$ 11.5 95.8 $\pm$ 9.04 0.60 $\pm$ 0.01
UDLM D-CFGγ=5 1013.6 $\pm$ 2.51 64.0 $\pm$ 5.1 0.62 $\pm$ 0.00

### v3

@@ 4.2 Improved Likelihood Bounds in Continuous Time
We now present our contribution to uniform noise discrete diffusion modeling. We leverage the above formulation to develop an improved NELBO by taking $T\rightarrow\infty$ and analyzing each term ${\mathcal{L}_{\text{recons}}},{\mathcal{L}_{\text{diffusion}}},{\mathcal{L}_{\text{prior}}}$ in (13). This yields three improvements: (1) a simple and elegant closed-form expression for the NELBO that is easier to reason about; (2) an analytical reduction of ${\mathcal{L}_{\text{recons}}},{\mathcal{L}_{\text{prior}}}$ to zero, which tightens the NELBO; (3) a further tightening via the continuous-time extension of ${\mathcal{L}_{\text{diffusion}}}$, as in Kingma et al. (2021) and Sahoo et al. (2024a; b).
@@
Reconstruction Loss (${\mathcal{L}_{\text{recons}}}$)  For ${\mathcal{L}_{\text{recons}}}$, if our noise schedule is such that $T\rightarrow\infty\implies\alpha_{t(0)}=1$ (i.e., $t(0)=1/T$), then the marginal $q({\mathbf{z}}_{\frac{1}{T}}\mid{\mathbf{x}})\rightarrow\mathrm{Cat}({\mathbf{z}}_{\frac{1}{T}};{\mathbf{x}})$. That is, in the limit, the first latent vector is identically equal to the clean data. We can thus parameterize our denoising network such that at time $t(0)$ the function simply copies its inputs: ${\mathbf{x}}_{\theta}({\mathbf{z}}_{\frac{1}{T}},1/T)={\mathbf{z}}_{\frac{1}{T}}.$ Additionally, we note that our choice of parameterization for $p_{\theta}$ implies that $p_{\theta}({\mathbf{x}}\mid{\mathbf{z}}_{{\frac{1}{T}}})={\mathbf{x}}_{\theta}({\mathbf{z}}_{{\frac{1}{T}}},1/T).$ Thus in the continuous time limit, we have:
$\lim_{T\rightarrow\infty}\mathbb{E}_{q}[{\mathcal{L}_{\text{recons}}}]=\lim_{T\rightarrow\infty}\mathbb{E}_{q}[\log p_{\theta}({\mathbf{x}}\mid{\mathbf{z}}_{\frac{1}{T}})]=\lim_{T\rightarrow\infty}\mathbb{E}_{q}[\log(\langle{\mathbf{x}},{\mathbf{x}}_{\theta}({\mathbf{z}}_{\frac{1}{T}},1/T)\rangle)]=0.$
Diffusion Loss (${\mathcal{L}_{\text{diffusion}}}$)  Turning finally to the diffusion loss term ${\mathcal{L}_{\text{diffusion}}}$, we first define the shorthand $D_{\mathrm{KL}}[q_{t}||p_{\theta}]=D_{\mathrm{KL}}[[q({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t},{\mathbf{x}})||p_{\theta}({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t})]]$ and then re-write this loss term as an expectation over $t$ uniformly sampled from $1/T,2/T,\ldots,T$:
@@
where ${\mathbf{x}}_{j}$ denotes the $j^{\text{th}}$ index of a vector ${\mathbf{x}}$, $\bar{\mathbf{x}}=N\alpha_{t}{\mathbf{x}}+(1-\alpha_{t}){\bm{1}}$, $\bar{\mathbf{x}}_{\theta}=N\alpha_{t}{\mathbf{x}}_{\theta}+(1-\alpha_{t}){\bm{1}}$, and we define $i=\arg\max_{j\in[N]}({\mathbf{z}}_{t})_{j}$ to be the non-zero entry of ${\mathbf{z}}_{t}$.
Combining our arguments regarding ${\mathcal{L}_{\text{prior}}}$ and ${\mathcal{L}_{\text{recons}}}$ with (17), yields our final tight bound:
@@ 5 Experiments
Datasets  For our language modeling experiments, we examine several discrete domains: reference genomes from tens species (Species10), the QM9 small molecule dataset (Ruddigkeit et al., 2012; Ramakrishnan et al., 2014), where molecules are represented by SMILES strings (Weininger, 1988), CIFAR10 discretized images (Krizhevsky et al., 2009), and three NLP datasets consisting of text8 (Mahoney, 2011), Amazon Review (McAuley & Leskovec, 2013; Zhang et al., 2015), and the one billion words dataset (LM1B; Chelba et al. (2014)). These datasets cover a range of domains and vocabularies of varying sizes (see Table 1). For guidance experiments, we explore species-specific sequence generation, molecular property maximization, and class-conditional image generation.
@@ 5.1 Language Modeling with Uniform Noise Discrete Diffusion
Our language modeling experiments show that (1) contrary to a widely-held belief, uniform noise diffusion can attain state-of-the-art performance on small vocabulary datasets (Table 1), and that (2) our UDLM are state-of-the-art among uniform noise diffusion (Tables 2 and 3).
Table 1: UDLM performs best with smaller vocabs. Perplexity ($\downarrow$) on various datasets. Best values are bolded. $**{}^{\text{*}}$ indicates values reported from early stopping on the validation set; otherwise validation performance at the end of training is used. From Sahoo et al. (2024a). \$From Lou et al. (2023).
||||Vocab.|||| AR MDLM UDLM
Species10 12 2.88 3.17 3.15
$**{}^{\text{*}}$ 40 2.19 2.12 2.02
CIFAR10 256 - 9.14 11.21
text8 35 2.35\$ 2.62 2.71
@@
In Table 1, despite previous evidence indicating that absorbing-state discrete diffusion greatly outperforms uniform noise, we find a more nuanced story. Namely, for smaller vocabulary regimes, the gap between MDLM and UDLM is negligible, with UDLM even outperforming absorbing-state on certain datasets. Even within a single domain, we find a trend between vocabulary size and performance gap, with the text8 results of UDLM on par with MDLM, and a more persistent gap for larger vocabularies. Intuitively, for larger vocabulary regimes, uniform noise diffusion models need to predict over a combinatorially larger set of potential clean data sequences compared to absorbing-state; smaller vocabularies reduce this complexity.
Figure 2: $T\rightarrow\infty$ improves validation PPL ($\downarrow$) on Amazon dataset.
Secondly, despite the persisting gap between absorbing and uniform noise discrete diffusion for larger vocabulary NLP datasets, we note that models trained with UDLM close this gap, as evidenced in Tables 2 and 3, where we show that UDLM attains the best reported uniform noise discrete diffusion language modeling performance.
Ablating Continuous Time Formulation  In Figure 2, we see the effect of increasing $T$ and using our continuous time NELBO. The curves represent validation perplexity on the Amazon Polarity dataset at various training steps, and we observe that increasing $T$ indeed improves language modeling with UDLM, i.e., $T=\infty$, performing best.
Table 2: UDLM outperforms other uniform discrete diffusion on text8. Best value is bolded & best uniform diffusion value is underlined. From Lou et al. (2023). $**{}^{\text{*}}$rom Shi et al. (2024).
@@
IAF/SCF (Ziegler & Rush, 2019) 1.88
Argmax Flow (Hoogeboom et al., 2021) 1.39
Discrete Flow (Tran et al., 2019) 1.23
@@
Mult. Diffusion (Hoogeboom et al., 2021) 1.72
MAC (Shih et al., 2022) 1.40
BFN (Graves et al., 2023) 1.41
D3PM Absorb (Austin et al., 2021) 1.45
SEDD Absorb (Lou et al., 2023) 1.39
MDLM (Sahoo et al., 2024a) (Retrained) 1.38
$**{}^{\text{*}}$ (Shi et al., 2024) 1.37
GenMD4 (Shi et al., 2024) 1.34
@@
D3PM Uniform (Austin et al., 2021) 1.61
SEDD Uniform (Lou et al., 2023) 1.47
@@
Table 3: UDLM outperforms other uniform discrete diffusion on LM1B. Best value is bolded & best discrete uniform diffusion value is underlined. From Sahoo et al. (2024a). $**{}^{\text{*}}$rom Lou et al. (2023). \$From Austin et al. (2021).
@@
BERT-Mouth (Wang & Cho, 2019) 142.89
D3PM Absorb (Austin et al., 2021) 77.50
Diffusion-LM (Li et al., 2022) 118.62
DiffusionBert (Wang & Cho, 2019) 63.78
SEDD Absorb (Lou et al., 2023) 32.79
MDLM (Sahoo et al., 2024a) 27.04
@@
D3PM ${}^{\text{\$}}$ (Austin et al., 2021) 137.9
SEDD $**{}^{\text{*}}$ (Lou et al., 2023) 40.25
@@ 5.2 Guided Discrete Diffusion
Our guidance results indicate that (1) classifier-free guidance is more useful when paired with diffusion models compared to AR (Table 4) and that (2) our proposed D-CBG is the best classifier-based method for discrete guidance, especially when combined with UDLM (Table 5 & Figure 3).
@@
Baselines  For guidance experiments, our primary baseline is the dominant AR approach. We compare to three flavors of guided AR. The first is applying D-CFG to AR models. We also use the established control mechanisms of Plug-and-play language models (PPLM; Dathathri et al. (2019)) and FUDGE (Yang & Klein, 2021). To demonstrate the better performance of our D-CBG method, we compare to Gruver et al. (2024) (NOS), an extension of PPLM to discrete diffusion.
@@
Hyperparameters  For both D-CFG and D-CBG, we vary the strength of the $\gamma$ parameter. Although the original FUDGE formulation simply uses $\gamma=1$, we also perform a search for this baseline. For PPLM and NOS, we vary the parameters of the Langevin sampling that updates models’ hidden representation, i.e., number of update steps $n$, step size $\eta$, and fluency KL weight $\gamma_{kl}$ (see Dathathri et al. (2019) and Gruver et al. (2024) for more details). For all methods, we display the best performing hyperparameter configuration in the main table results, and defer the the full sweep to Appendix E.
@@
Table 4: Diffusion decoding with D-CFG is more controllable than AR for genomic sequences. Mean $\pm$ standard deviation reported from five random seeds. Best values are bolded.
Model D-CFG $\gamma$ Disc. AUROC ($\downarrow$) F1 ($\uparrow$)
Random 1.00 0.07
AR 1 0.53±${}_{\pm\text{0.075}}$ 0.87±${}_{\pm\text{0.011}}$
AR 2 0.90±${}_{\pm\text{0.051}}$ 0.81±${}_{\pm\text{0.009}}$
AR 3 0.97±${}_{\pm\text{0.018}}$ 0.74±${}_{\pm\text{0.022}}$
MDLM 1 0.51±${}_{\pm\text{0.059}}$ 0.88±${}_{\pm\text{0.011}}$
MDLM 2 0.74±${}_{\pm\text{0.037}}$ 0.91±${}_{\pm\text{0.009}}$
MDLM 3 0.93±${}_{\pm\text{0.031}}$ 0.78±${}_{\pm\text{0.007}}$
UDLM 1 0.52±${}_{\pm\text{0.044}}$ 0.91±${}_{\pm\text{0.01}}$
UDLM 2 0.61±${}_{\pm\text{0.043}}$ 0.93±${}_{\pm\text{0.009}}$
UDLM 3 0.87±${}_{\pm\text{0.051}}$ 0.94±${}_{\pm\text{0.007}}$
Species-specific Genome Generation  For genomic sequences, we evaluate D-CFG with diffusion compared to with AR. We train on sequences of 32,768 nucleotides, using base-pair level tokenization and conditionally generate 64 sequences for each class. We train a small classifier to distinguish between generated and validation set sequences and report the area under the receiver operator curve for this classifier (Disc. AUROC). Values closer to 0.5 indicate that the classifier is unable to distinguish between synthetic and true sequences (Sarkar et al., 2024). To measure the controllability, we train a separate classifier on the Species10 dataset and measure macro F1 score of this ‘oracle’ classifier on the generated sequences. Results of this experiment are presented in Table 4. For reference, we also provide metrics for randomly generating sequences with nucleotide frequencies proportional to species representation in the data. We find that both MDLM and UDLM are able to better generate sequences that match the desired control parameter, with higher F1 scores relative to AR. Moreover, UDLM is able to outperform MDLM in satisfying this control. Importantly, we find that only UDLM is amenable to increasing the guidance parameter $\gamma$, where its metrics improves while AR and MDLM metrics degrade. Finally, of note, the diffusion model generation for this experiment is accomplished with far fewer function evaluations compared to AR. Whereas AR must decode each of the 32,768 tokens, because MDLM and UDLM can decode multiple tokens in parallel, we generate with $T=512.$
@@
Table 5: Guidance with discrete diffusion models better balances the generation of valid and novel molecules with maximizing the property of interest, drug likeness (QED) / ring count, compared to AR. Validity, novelty, and mean QED / ring count for novel sequences are measured for generated sequences from each method. Mean $\pm$ standard deviation reported from five random seeds. Best values are bolded.
QED Ring Count
Method Guidance Valid ($\uparrow$) Novel ($\uparrow$) Mean ($\uparrow$) Valid ($\uparrow$) Novel ($\uparrow$) Mean ($\uparrow$)
Original Data 133k 133k 0.47 133k 133k 1.74
@@
AR D-CFG 946.4±${}_{\pm\text{9.0}}$ 79.4±${}_{\pm\text{6.4}}$ 0.60±${}_{\pm\text{0.00}}$ 441.4±${}_{\pm\text{11.1}}$ 77.8±${}_{\pm\text{2.3}}$ 4.83±${}_{\pm\text{0.08}}$
MDLM D-CFG 317.4±${}_{\pm\text{11.5}}$ 95.8±${}_{\pm\text{9.0}}$ 0.60±${}_{\pm\text{0.01}}$ 90.0±${}_{\pm\text{8.2}}$ 60.0±${}_{\pm\text{7.5}}$ 5.26±${}_{\pm\text{0.15}}$
UDLM D-CFG 1013.6±${}_{\pm\text{2.5}}$ 64.0±${}_{\pm\text{5.1}}$ 0.62±${}_{\pm\text{0.00}}$ 998.2±${}_{\pm\text{4.5}}$ 216.2±${}_{\pm\text{13.0}}$ 4.88±${}_{\pm\text{0.04}}$